OpenStudy (anonymous):

rewrite this expression in terms of cosine cot(x)/2 csc(x)

OpenStudy (anonymous):

What is cscx

OpenStudy (anonymous):

Rewrite the problem properly

OpenStudy (anonymous):

Cotangent theta over 2 cosecant theta. I'm supposed to write that in terms of Cosine. I have it down to cos theta over sine theta times 1/2.

OpenStudy (mathstudent55):

\(\cot \theta = \dfrac{\cos \theta}{\sin \theta} \) \(\csc \theta = \dfrac{1}{\sin \theta} \)

OpenStudy (mathstudent55):

Write a complex fraction using what the cot and csc are equal to.

OpenStudy (anonymous):

Thanks you all! I figured it out before, but thanks again! This stuff is so fun to solve! -)